1. Prove that h(x) = x2 - x - 6  has a root on [ -4 , 0 ]
1
Expert's answer
2022-04-12T06:56:53-0400

ANSWER

First note, that the function is defined and continuous on "[-4,0]" . "h(-4)=(-4)^{2}-(-4)-6=16+4-6=14>0" , "h(0)=-6<0" . So "h(0)<0<h(-4)." By the Intermediate Value Theorem there exists a number "c\in(-4,0)" such that "h(c)=0" .

Note: the root of the equation is easy to calculate by solving the equation "x^2-x-6=0."

"c _{1}=\frac{1+\sqrt{1+24}}{2}= \frac{1+5}{2}=3 \notin( -4,0)", "c _{2}=\frac{1-\sqrt{1+24}}{2}= \frac{1-5}{2}=-2 \in (-4,0)" .

"c=c_{2}".